OpenStudy (anonymous):

anti derivative of : 2/(x^.5)

OpenStudy (anonymous):

u meant intergral?

OpenStudy (tkhunny):

Really big hint: \[\frac{d}{dx}\sqrt{x} = \frac{1}{2\sqrt{x}}\]

OpenStudy (raden):

anti derivative similar by integral go ahead @Otonashi, for helps

OpenStudy (anonymous):

then it is \[\frac{ -1 }{ 2x ^{4} }\]

OpenStudy (jusaquikie):

\[\frac{ 2 }{\sqrt{x} } =2x^{1/2} so F(x)=\frac{ 2x^{3/2} }{ 3/2} +C\]

OpenStudy (jusaquikie):

Just add 1 to your x and put it divided by your exponent, and don't forget your Constant

OpenStudy (tkhunny):

No. \[\frac{1}{\sqrt{x}} = x^{-1/2}\]

OpenStudy (jusaquikie):

that is the derivitave not the Anti-derivitave, this is the part of calculus 1 that leads into Integrals

OpenStudy (jusaquikie):

you are correct i made a mistake

OpenStudy (jusaquikie):

\[2x^{-1/2} = \frac{ 2x^{1/2} }{ 1/2 } +C = 4\sqrt{x}+C\]

OpenStudy (tkhunny):

Well, now you have an equivalence where you mean an antiderivative. Remember, notation means things. Don't write "=" unless you mean it.